OpenStudy (anonymous):

Find the angle between vector v and the positive x-axis (Give the absolute value of the angle found. Round your answer to two decimal places).

OpenStudy (anonymous):

The original vector is \[v=3i-9j\]

OpenStudy (loser66):



OpenStudy (anonymous):

Oh, okay. I think I see where to go next.

OpenStudy (loser66):

good

OpenStudy (anonymous):

So I am seeing: \[\cos (\theta) = \frac{ 9 }{ \sqrt{90}\sqrt{9} }\]

OpenStudy (loser66):

It seems good to me. just one more problem, wonder should we normalize those vectors first? or no need to do that. Ok, no need!! since we cancel out it then.

OpenStudy (loser66):

ok, final decision: take yours. It's good.

OpenStudy (loser66):

I am sorry for my opinions. I took the course 1year ago. Just make sure what I said.

OpenStudy (anonymous):

It's okay. Thank you so much for your help!

OpenStudy (loser66):

yw